### v1

# The EU AI Act, Stakeholder Needs, and Explainable AI: Aligning Regulatory Compliance in a Clinical Decision Support System

Abstract
Explainable AI (XAI) is a promising solution to ensure compliance with the EU AI Act, the first multi-national regulation for AI. XAI aims to enhance transparency and human oversight of AI systems, particularly “black-box models”, which are criticized as incomprehensible. However, the discourse around the main stakeholders in the AI Act and XAI appears disconnected. While XAI prioritizes the end user’s needs as the primary goal, the AI Act focuses on the obligations of the provider and deployer of the AI system. We aim to bridge this divide and provide guidance on how these two worlds are related. By fostering an interdisciplinary discussion in a cross-functional team with XAI, AI Act, legal, and requirements engineering experts, we walk through the steps necessary to analyze an AI-based clinical decision support system to clarify the end-user needs and assess AI Act applicability. By analyzing our justified understanding using an AI system under development as a case, we show that XAI techniques can fill a gap between stakeholder needs and the requirements of the AI Act. We look at the similarities and contrasts between the legal requirements and the needs of stakeholders. In doing so, we encourage researchers and practitioners from the XAI community to reflect on their role towards the AI Act by achieving a mutual understanding of the implications of XAI and the AI Act within different disciplines.
AI
artificial intelligence
ML
machine learning
CNN
convolutional neural network
DNN
deep neural network
ICU
intensive care unit
LSTM
long short-term memory
LOS
length of stay
NN
neural network
MTS
multivariate time series
RNN
recurrent neural network
XAI
explainable artificial intelligence
CDSS
clinical decision support system
1 Introduction
After long and complicated negotiations between the European Parliament, the EU Commission and the EU Member States, the AI Act entered into force in August 2024 (European Parliament and Council, 2024). AI-based systems are now regulated under a comprehensive set of product safety rules for the first time. In the coming years, different parts of the EU AI Act will become applicable. Providers and deployers of AI systems need to learn how to apply the different articles of the act, while notified bodies and certification authorities will need to establish a process to assess these systems according to the regulations. Many standardization efforts are already underway, but there will be significant uncertainty about how the AI Act has to be applied and how it relates to different aspects of AI currently under research.
Regulations such as the AI Act are especially important in high-risk artificial intelligence (AI) systems to ensure that they are safe and respect fundamental rights and values (Van Kolfschooten and Van Oirschot, 2024). The AI Act follows a risk-based approach: the higher the risk, the stricter the rule. Especially in high-risk systems, it describes many obligations for the provider and the deployer of AI systems.
[…]
cdss, which may fall under the definition of a high-risk AI system in the EU AI Act, place the responsibility for the high-risk decision on the end-user (typically, clinicians). However, the AI Act does not directly address end-user needs and instead primarily focuses on regulating the providers and deployers of AI systems. One of the few mentions of end-users is the statement that high-risk AI systems must be designed so that humans can effectively oversee these systems. But there are no obligations in terms of technical requirements that put the user’s needs at the center of its goals (Van Kolfschooten and Van Oirschot, 2024).
In contrast, nowadays more explainable artificial intelligence (XAI) researchers suggest putting end-users and their needs at the center of attention (Langer et al., 2021). XAI methods provide descriptions of the outputs of opaque AI models, thereby satisfying the needs of the end-users. This is very helpful in high-risk applications such as healthcare. Previous work has started exploring the role that XAI plays w.r.t. ensuring the trustworthiness of high-risk applications under the EU AI Act (Panigutti et al., 2023; Kiseleva et al., 2022; Longo et al., 2024; Esmaeilzadeh, 2024). XAI can contribute to implementing transparency and human oversight, which are key principles of the AI Act (Panigutti et al., 2023). Taking this line of thinking further, we believe that XAI—if desigend correctly—could play a crucial role in fulfilling the obligations of deployers and providers of AI systems. Therefore, we ask:
Research Question: How can Explainable AI bridge the gap between end-user needs and requirements of the AI Act?
To answer the research question, we analyze the relationship between the end-user’s XAI needs and the AI Act implications from a legal and technical perspective. We do this by using a concrete, AI-based clinical decision support system (CDSS) under development. This real-world application is currently under development in collaboration with clinicians and academic partners. The CDSS helps intensive care unit (ICU) coordinators who make choices about which patients can be moved out of the ICU with predicted key indicators such as the remaining length of stay (LOS). It is important to note that clinicians can use the predictions to make a decision in our CDSS — the system never decides on its own. To overcome acceptance and trust issues, we use XAI methods to provide human-understandable explanations of the system.
Research on the AI Act and research on XAI are often conducted independently, lacking interdisciplinary collaboration (Gyevnar et al., 2023). By providing a joint understanding of the role of XAI in the AI Act from an interdisciplinary perspective representing various stakeholder needs, we…
  1. 1.
    provide an analysis of the implications of the AI Act in relation to adopting AI in healthcare (see Section 4)
  2. 2.
    describe how the needs of the CDSS stakeholders can be captured through XAI (see Section 5)
  3. 3.
    assess the similarities and contrasts between the legal requirements of the AI Act and the stakeholder desiderata of XAI (see Section 6).
With this, our work is the first that analyzes the role of XAI in the AI Act in a cross-functional team, on a real-world CDSS that takes user-centered needs into account and contributes to the state-of-art by mapping obligations of the EU AI Act with the obligations of stakeholders and how to fulfill them using XAI. With this, we encourage researchers and practitioners from the XAI community to reflect on their role towards the AI Act by bridging the gap towards end-users.
2 Related Work and Background
In this section, we give an overview of the AI Act and XAI. We also discuss relevant related work. A number of studies already addressed the question of how AI can be applied in healthcare, using ethical and legal considerations (Aboy et al., 2024; Porsdam Mann et al., 2024; Karimian et al., 2022; Kiseleva et al., 2022). The function of XAI in fostering transparency and human oversight has also been the subject of considerable debate (Panigutti et al., 2023). Other works have dealt with the role of XAI in the AI Act (e.g., (Gyevnar et al., 2023)) and have given an overview of the AI Act and XAI (e.g., (Panigutti et al., 2023)). Nevertheless, it remains unexplored how XAI can be effectively utilized to fulfill the obligations of stakeholders as defined in the AI Act.
2.1 Background: The EU AI Act
In 2024 the European Union ratified the regulation of AI as technology (European Parliament and Council, 2024). The new regulation was motivated by a need to create trust for AI with respect to product safety, health and fundamental rights. The AI Act thus incorporates AI systems into the EU’s existing framework for product safety. Certain AI systems will need to undergo conformity assessment and subsequently carry the CE-mark as a sign of being a certified product, just like elevators, machinery and medical devices (see Section 2.1).
The regulation also gives authorities an important role in guaranteeing the responsible development and usage of AI, as public sector bodies are to conduct market surveillance but also foster AI innovation (see e.g., (Burden and Stenberg, 2022)). There are two central actors from the perspective of the regulation: the provider (see Section 2.1) and the deployer (see Section 2.1) of an AI system. These actors are obliged to take on responsibility for product safety.
CE-marking AI Systems
The conformity assessment for CE-marking AI systems can then be done in two ways. The first is through self-assessment where the provider of the system certifies that the requirements are met. This is also referred to as internal conformity assessment. The second conformity assessment procedure involves a notified body, an organization that has been accredited to be competent in conformity assessment of certain products. The role of the notified body is to give a neutral assessment to what extent the requirements are met. The nature of the AI system will determine which process is to be applied (Burden et al., 2024).
Obligations of the Provider
The provider is the organization creating the system, or as defined in the AI Act, Art. 3, a provider is “a natural or legal person, public authority, agency or other body that develops an AI system or a general-purpose AI model or that has an AI system or a general-purpose AI model developed and places it on the market or puts the AI system into service under its own name or trademark, whether for payment or free of charge”.
Providers of high-risk AI systems must adhere to several obligations as outlined in Art. 16-21. They must ensure compliance with requirements in Section 2 of the AI Act (Art. 9-15) and have a quality management system per Art. 17. They need to maintain documentation and keep logs generated by their high-risk AI systems. The systems must undergo a conformity assessment, have a CE-mark, and be registered in the Commission’s database. Providers also have to cooperate with authorities and take corrective measures if the system fails to comply with the AI Act.
The provider needs to perform risk management as outlined in Art. 9, which involves analyzing risks and developing mitigation strategies for the AI system. Transparency requirements in Art. 13 ensure that operations are clear enough for deployers to interpret the system’s output and use it appropriately. Human oversight, as specified in Art. 14, requires that the person operating the AI system understands the relevance of both input and output in the context of specific activities. Other technical requirements include data governance (Art. 10), technical documentation (Art. 11), record keeping (Art. 12), accuracy, robustness, cyber-security (Art. 15), and quality management (Art. 17).
There will be harmonized standards issued by the European standardization organizations CEN and CENELEC that further details the requirements in Art. 9 to 15 and 17. An overview of the standardization effort can be found on the two organizations’ joint technical committee’s website (European Committee for Standardization, 2025).
Obligations of the Deployer
According to the definition in Art. 3, a deployer is “a natural or legal person, public authority, agency or other body using an AI system under its authority except where the AI system is used in the course of a personal non-professional activity”.
The obligations for deployers of high-risk AI systems, as outlined in Art. 26, include taking technical and organizational measures to ensure systems are used according to instructions. If a risk is identified, deployers must promptly inform the provider and relevant authorities and suspend system use. Furthermore, deployers must assign human oversight to competent individuals with necessary training and authority. Deployers controlling input data must ensure it is relevant and representative in view of the intended purpose of the system. They are responsible for maintaining logs appropriately and informing individuals when decisions affecting them involve high-risk AI systems. Additionally, deployers must cooperate with competent authorities in actions related to the AI system to implement the regulation.
2.2 Background: XAI Terminology
As indicated, current research on explainable artificial intelligence (XAI) provides a new perspective on the relationship between an AI system and their affected parties. Therefore, we provide a brief nomenclature of relevant key concepts and terms. We define the key terms related to XAI, examine the various stakeholder terminologies, and clarify a common terminology of XAI methods.
Explainability, Interpretability, and Transparency
The terms explainability, interpretability, and transparency are often used interchangeably (Al-Ansari et al., 2024; Renftle et al., 2024; Hoffman et al., 2023; Barredo Arrieta et al., 2020; Mohseni et al., 2021). We distinguish these terms by categorizing them into approaches like Explainable AI and Interpretable AI, and desired properties such as Interpretability, Transparency, and Explainability (Mohseni et al., 2021). Explainable AI provides human-understandable explanations for predictions made by complex AI systems that are not inherently interpretable (Wang et al., 2019). Interpretable AI, on the other hand, requires no additional methods for human understanding, as these systems are inherently interpretable (Mohseni et al., 2021). Interpretability refers to the ability to explain or provide meaning in understandable terms, while Transparency indicates a model is interpretable by design (Barredo Arrieta et al., 2020).
Stakeholder Terminology
XAI goals such as trustworthiness, informativeness, and fairness have been widely studied (Al-Ansari et al., 2024; Renftle et al., 2024; Hoffman et al., 2023; Barredo Arrieta et al., 2020; Mohseni et al., 2021). Recently, these goals have shifted to focus more on specific stakeholders (Samek and Müller, 2019; Biecek and Samek, 2024), their desiderata (Langer et al., 2021), and expertise levels (Mohseni et al., 2021). Stakeholders’ interests, goals, expectations, needs, and demands regarding AI systems are referred to as their desiderata (Langer et al., 2021).
Understanding different stakeholders in XAI is essential for creating systems that meet various desiderata (Langer et al., 2021). Different groups have distinct interests, and recognizing these differences is key to developing methods that align with their specific desiderata. The success of an XAI method relies on satisfying stakeholders’ desiderata, which drives explainability approaches that enhance understanding. Langer et al. (2021), following Barredo Arrieta et al. (2020), categorize stakeholders into five classes: users, (system) developers, affected parties, deployers, and regulators. This classification ensures all relevant needs are considered, with the challenge being to provide suitable explanations for each group (Mohseni et al., 2021).
Following Mohseni et al. (2021), end-users are divided into three groups based on their expertise level: Novice Users, who use AI systems daily with little machine learning (ML) knowledge; Data Experts, including data scientists and domain experts using ML for analysis; and AI Experts, comprising ML scientists and engineers designing and implementing ML algorithms and XAI methods.
Taxonomy of Explainability Methods
We utilize common XAI taxonomies relevant to our CDSS system, following Speith (2022).
Stage refers to when the XAI method is applied. Ante-hoc methods are inherently interpretable AI models due to their simplicity. Post-hoc methods explain non-transparent ML models, like deep neural networks, using independent algorithms to provide feature relevance or visual explanations.
Applicability refers to post-hoc methods that are either model-specific, using the model’s internal structure, or model-agnostic, applicable to any ML model.
Scope refers to local explanations that detail specific input-output relations, while global explanations describe the overall ML model. Examples include model visualizations (Liu et al., 2017) and decision rules (Lakkaraju et al., 2016).
Other taxonomies include input data (e.g., image, time-series), output format (e.g., numerical, visual), and problem type (e.g., classification, regression).
2.3 Related Work
As previously mentioned, there are several studies that address the question of how AI can be applied in healthcare in general. Table 1 provides an overview of the related work papers and indicates how we position our own contribution.
Table 1: The related work can be divided into work that explores the role of XAI in the AI Act and work that analyzes the AI Act in healthcare from XAI or user perspectives. Our work extends these preceding studies by combining the user-centric view with the XAI view on the AI Act in a real-world healthcare scenario.
Paper AI Act XAI Healthcare Real World User-Centric
Gyevnar et al. (2023)
Pavlidis (2024)
Fresz et al. (2024)
Onitiu (2023)
Solaiman and Malik (2024)
Van Kolfschooten and Van Oirschot (2024)
Panigutti et al. (2023)
Our Work
Gyevnar et al. (2023) explore the transparency gap between the AI Act and XAI understanding, emphasizing the differing conceptualizations. They focus on stakeholder engagement and how XAI can address these concerns, aligning with our approach. Identified solutions include concentrating on legal definitions, AI system output, and data-related requirements to meet transparency standards. It’s important to note that transparency requirements are broad and should not be seen as concrete engineering instructions. XAI contributes to transparency but should not be viewed as a standalone solution, instead facilitating transparency while adhering to its foundational principles.
Pavlidis (2024) explores the role of XAI as a tool to bridge the gap between the opaque nature of ML decisions and the human demand for comprehensibility. In essence, XAI serves as a mechanism to align the objectives of the AI Act, such as enhancing transparency, with the preferences of end-users. The author examines the integration of XAI into the AI Act, with a focus on critical issues like standard setting, oversight, and enforcement. Fresz et al. (2024) review legal requirements for explainability of AI systems from the AI Act. Furthermore, they describe legal bases for the use of XAI and derive requirements the XAI methods need to fulfill. All three articles deal with the role of XAI in the AI Act. Our work extends this view by elaborating the role of XAI in a real-world healthcare scenario and by having a incorporating the the end-user needs.
A more domain-specific approach is presented by Onitiu (2023), who aims to translate technical safeguards in Art. 13-14 of the AI Act proposal for the use of medical diagnostic systems. The author proposes a broader perspective on the socio-legal implications of technical safeguards, such as post-hoc explainability for decision-making. The author emphasizes that post-hoc XAI methods must focus on harmonizing human expertise with patient-oriented values.
Similarly, Solaiman and Malik (2024) explore the AI Act in the healthcare domain with a user-centric focus on AI. Hereby, XAI plays a subordinate role. Instead, the authors focus on the doctor-patient relationship that evolved from a relation where the doctors held the dominant role and the patient follows their guidance without explanation, to a relationship where the patient is seen as active participants in their healthcare decisions. AI and especially XAI can have an important impact providing more information to both clinicians and patients, thus enhancing patient autonomy.
Van Kolfschooten and Van Oirschot (2024) analyze the implications of the AI Act for the healthcare domain. Here, they give an overview of the obligations for various healthcare stakeholders. Then they use their insights to provide a guideline to address the unique needs of the stakeholders and make recommendations for the implementation and standardization of the act. They also highlight the importance of collaborating with different stakeholders like AI developers, healthcare professionals, patient communities, and regulators to develop those guidelines.
The preceding three studies have all examined the AI Act in the context of healthcare from the perspective of users and XAI. In contrast to these studies, our approach involves a combination of both user-centric and XAI perspectives, which entails the analysis of healthcare stakeholders from the perspectives of AI Act and XAI.
In a more application-oriented study, Panigutti et al. (2023) conduct an interdisciplinary analysis of the AI Act’s approach to opaque AI systems, aiming to clarify XAI’s role. They focus on Art. 13, concerning transparency, and Art. 14, regarding human oversight, providing an overview of XAI terminology and key concepts. The authors also review the XAI field, noting challenges and limitations consistent with previous findings. They analyze how the AI Act addresses AI system opacity through transparency and human oversight, presenting a use case involving an AI-based proctoring system without XAI tools. In contrast, our study extends beyond Art. 13 and 14 and examines other articles that establish obligations for healthcare stakeholders.
3 Methodology
To answer our research question, we apply a two-pronged methodology within a cross-functional team. This section first describes the team and the distribution of responsibilities and incentives. We then describe the CDSS. That is followed by the two prongs of the approach, first the legal analysis of the AI Act and then the XAI analysis, both in relation to the CDSS. We then show how the two prongs merge.
3.1 The Cross-functional Team
The team behind this contribution consists of five individuals. Two individuals represent a company developing the CDSS. One of them has a PhD in computer science and has extensive experience in requirements engineering and empirical software engineering. The other is a PhD student who focuses on the human-centered implementation of XAI methods for AI-based CDSSs. Human-centered XAI methods put the human and their desiderata at the center of the implementation of XAI methods. The daily work of these individuals takes place in the context of an interdisciplinary exchange with the clinicians in a publicly funded research project. They organized the workshops (see Section 4) with usability engineers from the company, but are not part of the stakeholders that have been analyzed.
Two individuals represent a research institute; one with a background as a judge at an appeals court as well as working for multiple governmental investigations regarding the introduction of new laws, the other with a PhD in computer science with a focus on model-driven engineering and natural language processing. Together they have assisted the Government’s Office during negotiation and implementation of the AI Act and have shared their insights with the EU Commission. They also assist a notified body in defining the implications of the AI Act to existing and possible new business opportunities, e.g., when is a notified body needed, how to show fulfillment within a conformity assessment procedure and to which concrete products is the AI Act applicable.
The fifth member of the author team is a university professor for information systems and AI. The main responsibility in relation to this contribution is as an external auditor to assess the validity and scientific soundness of the contribution and deliver both a technical and socio-technical viewpoint.
3.2 The Clinical Decision Support System
The AI-based CDSS is designed to help ICU coordinators manage bed occupancy. When the ICU reaches capacity and new patients require intensive care, coordinators must decide which patients can be moved to another ward and which must remain in the ICU. Typically, such decisions rely on extensive clinical experience and a clinician’s intuitive sense of whether a patient is stable enough to be transferred. However, when these decisions fall to assistant doctors who lack the same depth of experience, the quality of the decision-making process may be compromised.
This is where the CDSS comes into play. It supports clinicians by predicting key indicators that guide the transfer decision—such as remaining LOS, mortality probability, and readmission probability—using vital signs like heart rate and oxygen saturation from the past 24 hours. In particular, when the AI’s predictions conflict with a clinician’s intuition, the interpretability of the system is crucial for reconsidering and validating the final decision. To generate these predictions, a DNN is trained on time-series patient data spanning the years 2017 to 2024. Notably, the system does not suggest which patients to move, but only shows the predictions to help clinicians make decisions themselves.
3.3 Legal Analysis
The legal analysis consists of mapping the different articles of the AI Act to the system under investigation – which technical and organizational requirements does the system imply? Then, who will be responsible for the fulfillment of the requirements? The details of the technical requirements are not as important for the analysis as their implications for the different roles. For the legal analysis, e.g., it is more important to know who is responsible for which aspects of data management (input data, validation data, etc.), than how to assess the relevance or completeness of data. As analyses emerge we visualize the concepts and their relationships since text has the drawback of enforcing a linear representation (see e.g. Figure 2(a)). The primary outcome of the legal analysis is given in Section 4.
At the same time we make decisions on what is in and out of scope. The latter is necessary since the AI Act consists of 113 articles that define roles, responsibilities, technical requirements, fines and so on. That is complemented by 13 annexes that supply further detail to the articles. To motivate and help with interpretation of the articles and the annexes, there are 180 recitals. Knowing what to include in an analysis is therefore an analysis in itself and something that is made easier as more is known about the system under consideration. In the case of this contribution we settle on focusing primarily on the responsibilities of the roles of the provider and deployer, leaving out the content related to what happens if a system does not comply. We also keep the level of detail abstract in terms of technical requirements as going into detail would yield enough content for multiple academic contributions per article of the AI Act.
In parallel we conduct analyses of the AI Act in relation to the new machinery regulation as an assignment for the notified body (c) and a first analysis of the cyber-security requirements of the AI Act in relation to the regulation on medical devices and the Cyber Resilience Act (anonymized). These parallel investigations also suffice to better understand the AI Act in relations to the CDSS.
3.4 Requirements Engineering Activities
To understand the requirements of the clinicians that are going to use the CDSS, we conduct a workshop in which we collect the most important functional and quality requirements. The workshop is modeled after the Design Thinking method (Ferreira Martins et al., 2019), but much abbreviated. Participants were clinicians and ML researchers, with a requirements engineer moderating the workshop.
We engage in brainstorming sessions in which we map out the processes in the ICU and identify the most relevant user stories. These are then transformed into a user journey using user journey mapping (Milicic et al., 2012) to capture the requirements on the system implied by these processes. A scoping exercise then allows us to narrow down which aspects of the process can be supported by the system and thus define the scope of the CDSS.
To refine the requirements and the desiderata for the XAI methods in particular, we conduct several interviews with clinicians. We also perform contextual inquiry (Bednar and Welch, 2009) where we follow a clinician on their rounds in the ICU to better understand how the process we elicit works in practice and which factors they consider before deciding which patient to move out of the ICU.
3.5 Stakeholder Mapping and XAI Analysis
The next step of analysis is to identify the relevant stakeholder groups. Thereby, we have discussions involving the AI Act and XAI experts, following the stakeholder mapping approach outlined by Langer et al. (2021). Stakeholders identified during the Design Thinking Workshop are mapped, and we summarize their specific desiderata and expand upon them in regular project meetings and exchanges. For the choice of XAI methods, the categorization by Speith (2022) is used. The Ph.D. student analyzes the problem and selects common methods that align with the problem types and relevant requirements of the CDSS, as identified in the requirements engineering workshops.
3.6 The Merging of the two Approaches
The legal and the XAI analysis are made in parallel so that they benefit from insights from the other approach. This takes place over three different phases. Phase one was a one-day sprint (March 2024) where the AI Act is introduced to the company and a subsequent workshop starts to map out the high-level applicability towards the CDSS. Phase two occurs six months later (September 2024) to share new insights in a second workshop. The third phase runs from December 2024 to January 2025 and is conducted as six weekly meetings with the ambition to document the combined analysis and the reasoning. Between the meetings of the third phase, each individual contributes to the shared document and the meetings discuss the combined effort and changes are made as new insights were gained. The outcome in this contribution represents the final and polished analyses.
3.7 Validity Strategies
Following Lincoln and Guba (1985), we structure the validity strategies along Credibility (establishing an analysis true to the data and context); Transferability (ensuring that the findings were applicable beyond the context); Dependability (enabling repetition of our contribution); and Confirmability (limiting the impact of our own incentives and bias).
As seen in Section 3.6, we apply a methodology that allows for triangulation of methodologies applied separately by the individuals of the team, over time (Credibility and Confirmability). We also show how our analysis goes beyond a specific AI system in sections 6 and 7 (Transferability). The ambition is that it is possible to replicate the study as we describe the team behind the contribution and its incentives, as well as the applied methodologies for collecting and analyzing data (Dependability and Confirmability).
4 Applicability of the AI Act to the CDSS
We analyzed the CDSS in the context of the EU AI Act and establish whether the act is applicable to the case of the CDSS system. To do this, we need to answer three questions:
  • Does the system fall within the legal scope? (Art. 2)
  • Does the system fit the legal definition of an AI system? (Art. 3 (1))
  • Does the system need to be CE-marked? (Art. 6.)
Scope: The AI Act has a number of clauses defining the scope of the regulation. For instance, it “does not apply to any research, testing or development activity regarding AI systems or AI models prior to their being placed on the market or put into service” or systems used in a purely personal, non-professional activity. After going through the scope as defined in Art. 2, our assessment is that the AI system under investigation is within scope of the AI Act as it is to be put into use in a clinical context.
AI system: An AI system in the AI Act is defined as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments”.
A key concept here is infer which can be done by machine-learning approaches but also through knowledge- or logic-based technologies. However, “systems that are based on the rules defined solely by natural persons to automatically execute operations” should not be seen as AI systems (recital 12). Our assessment is that the AI system under investigation is within scope of the definition of AI system as it, from received input, generates a prediction that, if the clinician approves, will influence where and by who the patient will be treated.
CE-mark: There are two scenarios where the AI system needs to be CE-marked: the first covers clinical systems that fall under EU regulation 2017/745 on medical devices and subsequently need to undergo a conformity assessment procedure involving a notified body (Art. 6.1 of the AI Act); the second is for AI systems that are listed in Annex III of the AI Act. In the latter case, point 5.a refers to “AI systems intended […] to evaluate the eligibility of natural persons for […] healthcare services, as well as to grant, reduce, revoke, or reclaim such […] services”.
Our assessment is that, while it is unclear if the CDSS is a medical device or not, it does not fulfill the criteria to fall into a risk category mandating the involvement of a notified body, according to regulation 2017/745 on medical devices. However, in preparing for the regulated case scenario, we do assess that the AI system under investigation is covered by the list in Annex III. Some systems under Annex III can be exempt from CE-marking if they are solely used for preparatory purposes (Art. 6.3a.d), but since our system conducts profiling as defined in General Data Protection Regulation (GDPR), we cannot apply this exempt (Art. 6.3).
Taken together, this implies that the AI system will need to be CE-marked through internal compliance, i.e., without involving a notified body (Annex IV of the AI Act).
Conclusion: The AI system needs to be CE-marked through an internal conformity assessment procedure in accordance with the AI Act. This is the obligation of the provider. In practice, this means that the provider needs to fulfill several obligations, and described in Section 2.1. This includes fulfilling technical criteria (Art. 9-15) and having a quality management system (Art. 17).
That is mirrored by the obligations of the deployer (Art. 26), also described in Section 2.1. Since the system will be deployed by a legal person governed by public law (or by private entities providing public services), there is another obligation: a Fundamental Rights Impact Assessment shall be conducted by the deployer (Art. 27). How to do this exactly will be further detailed in a harmonized standard.
5 Choosing the right XAI Method
Choosing the right XAI method is important to ensure that the stakeholder needs are satisfied (Schoonderwoerd et al., 2021). Therefore, we analyze relevant stakeholders and elaborate their desiderata (see Section 5.1). This information helps us make well-founded choices about the XAI method in the implementation phase of the CDSS (see Section 5.2).
5.1 Our Stakeholders and their Desiderata
Table 2: Overview of the real-world stakeholders and desiderata mapping.
XAI Stakeholder Roles AI Act Stakeholder Roles Clinical Role Desiderata
(System) Developer Provider ML Researcher and Engineers Reliability, Performance
Deployer Deployer Hospital Management Acceptance, Legal Compliance
User Natural Person Clinicians Interpretability, Usability,
Trust/ Appropriate Reliance
Affected Parties Natural Person(s) Patients Correctness, Fairness
Regulator AI Office, AI Board, Competent Authorities, … Competent authorities Compliance
The first group of stakeholders are researchers and engineers specializing in ML, who are categorized as developers and identified as providers under the provisions of the AI Act. Their primary requirement for the AI system is the achievement of high performance and reliability. In order to fulfill this requirement, they concentrate on quality measures such as accuracy and robustness, which are of the highest importance for the practical implementation and success of the model in real-world applications. This will increase end-users’ acceptance of our AI system which is a key factor for hospital management.
As the deployer of the AI system, the hospital management is interested in acceptance of the AI system by the clinicians. It needs to be a reliable source of support in high-stakes decision-making processes. Furthermore, the AI system must comply with a range of regulations, such as the AI Act, to ensure its appropriate in the hospital.
Clinicians are the primary users of the AI system in CDSS, as they are the only individuals who directly interact with the AI outputs and ultimately make the final decisions. Most of these clinicians are novice users; as defined in Section 2.2, the AI system must be accessible and user-friendly. This translates into the desideratum interpretability, which involves providing comprehensible, plausible, and correct explanations for the AI system. This interpretability allows clinicians to question AI predictions rather than simply accepting them, thereby fostering trust and appropriate reliance (Schemmer et al., 2023). Clinicians should trust the AI system when it is correct and appropriately rely on their judgment when the AI system is incorrect. This is most important when the AI system makes a false prediction that negatively impacts the patient, e.g., the AI system falsely predicts a short remaining LOS for the patients, although the patient is very sick and will have to stay longer. In this situation, the clinicians must rely on their medical expertise.
Patients, although not directly interacting with the AI system, are significantly impacted by the decisions made by clinicians using the AI system and thus an “affected party” according to the definition in Section 2.2. Patients have a strong need for correct and fair decision-making. Fairness in AI-driven decisions is a critical requirement, as it directly affects patient outcomes and trust in the healthcare system.
The regulator corresponds in our case to the competent authority, which can be mapped to different AI Act stakeholder roles, e.g.: AI Office or AI Board.
5.2 Our XAI Method Choices
After we clarified the relevant stakeholders and their desiderata, we now describe the concrete XAI method choices we made for our CDSS. We made choices in the categories applicability, scope, and stage, as proposed by Speith (2022). Categories like input data and problem type are pre-determined by the problem the application addresses.
Input Data, Problem, and Model Type: The AI system is based on a complex DNN that processes multivariate time series data, including patient vital signs, to predict the remaining LOS of an individual patient.
Stage: While DNNs provide great performance by learning complex relationships in data, they are not intrinsically interpretable and lack transparency (Barnes and Hutson, 2024). Therefore, we employ post-hoc XAI methods to provide insights into this black-box model.
Applicability: To ensure a reliant and permanent AI model, we dynamically retrain and update the AI model to avoid data drifts. Some changes in the AI model might affect the AI architecture. Therefore, the model-specific XAI method could not be applied anymore. For this reason, we selected model-agnostic XAI methods. This choice allows us to apply the XAI method to any underlying AI model that is based on a neural network structure.
Scope: The primary goal of our XAI method is to satisfy our end user’s desiderata and thus target interpretability, usability, and trust or appropriate reliance. For this reason, our XAI method focuses on a local scope, enabling clinicians to understand specific predictions. This local scope is done by interpreting relationships between input data, such as patient vital signs, and predicted outcomes, like the estimated remaining LOS. This model’s interpretability is crucial to ensure an appropriate reliance on clinical decision-making.
Output Format and Result: To be sure that the output of the XAI method is understood accordingly, we are using a combination of different output formats for the XAI method result, such as visual and numerical information. This degree of freedom allows us to provide an appropriate output format for users with different desiderata and levels of expertise.
Figure 1: Example Heatmap Explanation with Shap. Shows the five most important features in descending order of relevance over a period of 14 hours. The color gradient indicates whether the feature value at a given timestamp has a positive (red) or negative (positive) impact on the predicted LOS.
Finally, the different choices in the categories allow us to choose between XAI methods such as Shap (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016), Integrated Gradient (Sundararajan et al., 2017), or Saliency Maps (Simonyan et al., 2014). Functionally-grounded evaluation methods (Doshi-Velez and Kim, 2017) that measure explanation properties such as complexity, faithfulness, or sensitivity help us to select the best XAI method (Chen et al., 2024).
For example, Figure 1 shows a concrete heatmap explanation highlighting the importance of features for all time series steps. Here, the clinician gets an overview of the five most important features over the last 15 hours of the patient in the ICU. The color gradient indicates if the specific feature value at the specific time stamp has a negative (green) or positive (red) effect on the predicted remaining LOS. A short LOS indicates a more positive state of health then a long LOS. This explanation of the LOS prediction assists the clinician in appropriately deciding whether the patient can be moved to another ward.
6 Contrasting XAI with the AI Act
From analyzing the same CDSS from the perspectives of XAI and the AI Act, we have created opportunities to reflect on what one perspective reveals about the other, thus a way to answer our research question.
The AI Act refers to stakeholder obligations, while XAI describes stakeholder desiderata. Mapping these two perspectives is not a straightforward one-to-one mapping. Rather than exhaustively detailing the inter-relationships, we explored a thin slice of their intersection to show how XAI can fulfill the AI Act’s legal requirements. This means we are explicitly expecting further research to complement and refute our findings (see KI1-KI3), both in scope and in conclusions if XAI is chosen as a relevant way to comply with the requirements of the AI Act. A schematic view of the analysis is found in diagrams of Figure 2.
6.1 Roles, Responsibilities and Needs
(a) The stakeholder roles and responsibilities when combining the AI Act with XAI.
(b) Introducing XAI to comply with legal obligations will involve the entire stack of stakeholders.
Figure 2: Stakeholder roles and their responsibilities in XAI and the AI Act. Italics denote XAI roles, bold denotes AI Act roles. Parenthesized roles are from the CDSS. An arrow from A to B denotes that A has a responsibility towards B.
As laid out in Section 2.2, the XAI community is concerned with stakeholder desiderata (Langer et al., 2021) while the AI Act concerns stakeholder responsibilities (see Section 2.1). While there is an overlap in the main stakeholder roles introduced by XAI and the AI Act, there are also differences. Where the XAI community is satisfied by the role of regulator (Langer et al., 2021), the AI Act has multiple roles. There is the national authority assigned to conduct market surveillance of AI systems, the notified body that inspects to what extent an AI system and the developer’s organization conform to the requirements and the EU Commission who all have a say in how the AI Act is to be implemented and enforced. Even a court could serve as a regulator if the system developer and the national authority have different opinions. In the case of our CDSS, the system developer can even take the XAI role of regulator. This can seem counter-intuitive, but it is one way of interpreting the conformity procedure based on self-assessment.
On the other side, the XAI community pays explicit attention to the end user and the affected parties (Langer et al., 2021), while the AI Act trusts the provider and the deployer to take care of those perspectives. As such the stakeholder roles can be seen as lacking a common ground, or to complement each other. That leaves ample room for the XAI community to define their stakeholder roles in relation to the responsibilities enforced by the AI Act.
Another observation is that while Langer et al. (2021) treat individuals “as representative members of specified stakeholder classes”, our analysis mandates us to distinguish between affected parties as concrete individuals and as an abstract collective due to the responsibilities of the deployer. Art. 26 in the AI Act defines requirements on the usage of the AI system on an individual level at run-time, while Art. 27 poses requirements on performing risk assessments of the AI system in relation to its intended usage pre-launch in respect to the affected parties in general. This is visualized in Figure 2(a) as two different responsibilities (arrows) between the deployer and the affected parties.
Key Insight (KI1) The XAI community should define their roles in relation to the responsibilities enforced by the AI Act.
6.2 Bridging the AI Act’s Gap
The AI Act deliberately excludes individual patients’ will, focusing instead on roles relevant to product safety. Interestingly, the XAI perspective does not give the patient any more agency than the AI Act does. Thus, both the AI Act and XAI have roles that act on the behalf of the affected party, which by itself lacks the ability, mandate, or obligation to change the functionality of the AI system. However, XAI emphasizes the needs of affected parties in relation to the user, potentially bridging the gap between end users and affected parties, which the AI Act perspective lacks (highlighted as blue arrow in Figure 2(a)).
Notably, the AI Act does not explicitly define stakeholder roles, instead referring to “relevant stakeholders”, “external stakeholders”, “multi-stakeholder” etc. It is therefore our analysis that defines Deployer, Competent Authority and Provider as stakeholder roles. Regarding the equivalence to the XAI roles of User and Affected parties, the AI Act refers to the former as “staff” (the collective that shall be informed, trained and have responsibilities regarding AI systems according to Art. 17 on quality management systems), “affected workers” (those who will use the CE-marked system in relation to Art. 26 and the impact on individuals’ rights) and “natural persons” (those who have the adequate training for using a CE-marked AI system, Art. 26). The latter role is referred to as “categories of natural persons or groups of persons” (impacted by CE-marked systems under Art. 27). As the AI Act lacks systematic terminology for stakeholder roles, we use Natural person and Natural Person(s) to emphasize human involvement (natural person is legal terminology for human) rather than legal persons (legal terminology for organizations, companies, authorities, etc.).
Key Insight (KI2) XAI enhances the AI Act by focusing on the relation between the affected parties and the user and their needs.
6.3 From Regulator to Affected Parties—and back
The AI Act mandates respecting the fundamental rights of affected parties but does not specify technical methods or thresholds. This opens a research trajectory for the XAI community in how to achieve appropriate methods and reasonable cut-off points. As an example, how certain does a recommendation system need to be about key-indicators, such as the LOS, that support the decision which patient is moved from the ward to make room for another patient? How do we calculate that certainty and how is it represented to physicians, hospital management, regulators and affected parties? For this to be relevant in a legal context, the future contributions of the XAI community need to be understood and accepted by the diverse group of actors that can take the stakeholder roles of deployer, developer and regulator (see Figure 2(b)). This aligns with Pavlidis (2024), who emphasizes that XAI must operate in compliance settings to meet regulatory requirements. In our case, it led us to target interpretability, usability and trust as key needs to address.
The AI Act introduces operational activities like a data lifecycle and risk management aspects (e.g., the delegation of responsibilities), without specifying methods. Here the XAI community has an opportunity to delimit what is within its interests to contribute towards, but also new aspects of XAI development and usage for future contributions.
In our case, FRIA must be conducted before system deployment, requiring the deployer to assess the impact and risks to patients’ fundamental rights (Art. 27, which relates to Fairness). This links back to the provider’s obligations. The FRIA uses information from the AI-system provider (Art. 13) and the provider’s risk assessment (Art. 9). Technical documentation and logs help the provider conduct the FRIA and assist the user in evaluating output relevance, impacting XAI desiderata like trust and acceptance.
Key Insight (KI3) XAI facilitates compliance with the AI Act by offering essential documentation for FRIA, thereby enhancing trust and acceptance among end-users.
7 Discussion and future research
From the XAI perspective, the starting point of our analysis is the obligations of the deployer. The persons assigned the oversight or usage of the system need to have the right competence, authority, and support so that they can assess if the output is reasonable given the input and current state of the ICU (Art. 26 of the AI Act, which relates to the XAI term of Interpretability). As described in Section 3.2, the CDSS provides key indicator recommendations to support the clinical decision, which patients can be moved to another ward. While XAI can provide measures for explaining how the recommendation came to be, it does not assess whether it is right or wrong. Operating the ward still requires human judgment to assess which recommendations to follow and when alternative routes are more appropriate (see, e.g., (Schoeffer et al., 2024; Gomez et al., 2025) on Human-AI Collaboration).
The design choices we propose for the AI system under consideration are influenced by the stakeholders’ needs but also help us to fulfill the obligations of the AI Act, e.g.:
  • Risk Management: We employ local XAI methods such as SHAP that have the potential to enable clinicians to critically evaluate the predictions by the AI system. This is one way to mitigate the potential risk of the end-user relying uncritically on the AI. Therefore, XAI provides a risk mitigation measure as suggested by the AI Act in Art. 9 (2d).
  • Transparency: Our XAI method provides information that can help clinicians ensure an appropriate reliance on AI systems, thereby “enab[ling] deployers to interpret the system’s output and use it appropriately” (see Art. 13). Especially, finding the “appropriate type and degree of transparency” needs to involve stakeholders and take their individual desiderata into account. Transparency can thereby be seen as a precondition to XAI.
  • Human Oversight: An appropriate and individualized XAI method, in our case including complex outputs like heatmaps, is crucial for ensuring that high-risk AI systems can be effectively monitored by humans during their operational period, as stated in the AI Act (see Art. 14).
The overall ambition with the AI Act is to create trust in AI as a technology. It does so by stating general requirements that need to be addressed concretely, both in relation to the intended purpose of the system but also to the involved actors. Here XAI can be of use. As detailed in Section 5.1, the selection of XAI method must be tailored to the stakeholders and their desiderata. The success of this approach depends on meeting the specific desiderata and satisfaction of each end-user. Engaging relevant stakeholders in the selection of XAI method allows addressing their individual desiderata and an effective implementation of different elements of the AI Act.
Yet, there remains much to do. Since the XAI method choices reflect what happens between physician and patient (novice user and affected party), they need to be percolated upwards to the hospital management (deployer) of the AI system and captured as requirements, consequently fulfilled by the system developer (developer) of the AI system. Finally, the design choices and their implementation need to be understood and accepted by the regulator (see Figure 2(b)). XAI can serve as an instrument to build trust at all levels in relation not only to AI as technology but to a specific AI system – the regulator trusts the AI system to be fit for purpose, the deployer and provider trust the AI system to be fit for purpose and the user and the affected parties trust the AI system to be fit for purpose. A similar reasoning is applicable for the usability of the AI system — it needs to be usable for both the end user and the deployer. We foresee a number of studies to further explore the relationship between the levels but also between XAI as a tool and the legal requirements of the AI Act and how they are further detailed in the upcoming harmonized standards.
Here it is worth pointing to an emerging discussion within the academic community on the risks of introducing AI in the healthcare system, specifically in the relationship between clinician and patient and the liability concerns in relation to perceived mistreatment (see, e.g., (Solaiman and Malik, 2024; Duffourc and Gerke, 2023; Čartolovni et al., 2023)). XAI could both serve as a means to lessen the risk of litigation, but also in itself poses a risk if not implemented properly. This is exemplified by Weller (2022), who states that explanations can lead to unsubstantiated trust to the lay user, and hence have a negative impact on affected parties. If XAI is to realize its potential in relation to the AI Act it is important that it is not applied to “XAI-wash” unsuitable systems (see e.g., Rudin (2019) who criticizes the use of XAI methods that may not be suitable in high stakes decisions which are nonetheless justified using XAI), but for a genuine wish to make AI a suitable technology for human needs.
8 Conclusion
As the AI Act enters into force, AI-based systems are now regulated under comprehensive product safety rules. Therefore, the AI Act refers to stakeholder obligations without putting the end-users and their needs at the center of attention. On the other hand, the research around XAI focuses on stakeholder desiderata. By analyzing a CDSS under development in relation to the AI Act and by analyzing the stakeholders from different perspectives, we map stakeholders’ desiderata to the requirements of the AI Act. In doing so, we discuss the role of XAI in the AI Act and make the stakeholder role of the (novice) user and the subsequent relation to the affected parties explicit.
Especially in high-risk areas like healthcare, we emphasize the critical role of human oversight and competence in deploying AI systems to ensure trustworthy outputs and appropriate decisions. We highlight the necessity of conducting FRIA and engaging stakeholders in selecting tailored XAI methods to build trust and acceptance. We see XAI as a chance to initiate a dialogue on the effective implementation of different elements of the AI Act and as an instrument to build trust in a specific AI system at all stakeholder levels.